1. Suppose you are provided with the following data on the components of aggregate expenditure of Ghana in the year 2020. Consumption function (C) = Gh¢2000m + 0.65Yd Tax function (T) = Gh¢1400m + 0.2Y Gross private domestic investment (I) = Gh¢1800m Government expenditure (G) = Gh¢1970 Exports (X) = Gh¢1300m Imports function (M) = Gh¢1220m + 0.2Y Use the following data above to compute the following: a) Marginal propensity to consume (MPC) out of national income b) Marginal propensity to save (MPS) out of national income c) The savings function as a function of national income d) What is the value of the equilibrium level of output: i. Using the aggregate expenditure approach ii. Using the injections leakages approach. e) What is the value of autonomous spending multiplier? f) What is the fiscal stance (the budget balance) of the government at equilibrium GDP? g) What is the nature of the trade balance at equilibrium GDP


1
Expert's answer
2021-07-07T09:03:13-0400

Given :

C = 2000 + 0.65 Yd

T = 1400 + 0.2 Y

I = 1800

G = 1970

X = 1300

M = 1220 + 0.2 Y


(a) Marginal propensity to consume = 0.65

Because Consumption = Autonomous consumption + Marginal propensity to consume * Income

(b) Marginal propensity to save = 0.35

Because Marginal propensity to consume + Marginal propensity to save = 1

So,

0.65 + MPS = 1

MPS = 0.35

(c) Savings function :

Income is either consumed or saved.

ie. Y = C + S

Y = 2000 + 0.65 Yd + S

S = -2000 + 0.35Yd

Therefore, Saving function is : S = - 2000 + 0.35Yd

(d) Equilibrium level of output :

(i) By aggregate expenditure approach, equilibrium is attained where,

Y = C + I + G + X - M

Y = 2000 + 0.65Yd + 1800 + 1970 + 1300 - 1220 - 0.2Y

Y = 5850 + 0.65 (Y - 1400 - 0.2Y ) - 0.2Y

Y = 5850 + 0.65Y - 910 - 0.13Y - 0.2Y

Y - 0.65Y + 0.13Y + 0.2Y = 4940

0.68 Y = 4940

Y = 7264.70


(ii) By injection- leakage approach

Equilibrium is attained where

S + T + M = I + G + X

-2000 + 0.35Yd + 1400 + 0.2Y + 1220 + 0.2Y = 1800 + 1970 + 1300

-2000 + 0.35 (Y - 1400 - 0.2Y) + 1400 + 0.2Y + 1220 + 0.2Y = 5070

-2000 + 0.35Y - 490 - 0.07 Y + 1400 + 0.2Y + 1220 + 0.2Y = 5070

0.68Y = 4940

Y = 7264.70

Therefore, equilibrium level of output is 7264.70